a rectangle is twice as long as it is wide. the perimeter is 72cm. what is the area of the rectangle
Answer by rajagopalan(174)

Let width=x
Length=2x
Perimeter=2(x+2x)=6x
we are given that 6x=72
so x=72/6=12
width=12
Length=2x=2*12=24 CM
Area=L*B=24*12=288 Sq. CM